a boat has travelled 25km in 7 hours 45 min.calculate speed of this boat in knots
Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

rate = distance / time = 25 / 7.75 = 3.23 km / hour
:
Note 1 km / hour = 0.54 knot
:
3.23 / 0.54 = 5.98
:
*******************************
Speed of boat is 5.98 knots
*******************************
:

Answer by MathTherapy(10552)


a boat has travelled 25km in 7 hours 45 min.calculate speed of this boat in knots